# Representing rational integers by generalized quadratic forms over quadratic fields

### 11 Mar 2024 v1

@@ 4. Generalized quadratic forms
If $G$ is a generalized quadratic form in $n$ variables over $K$, then we can define a quadratic form $Q$ in $2n$ variables by
@@
In other words, $Q$ is obtained from $G$ by expressing the variables $z_{1},z_{2},\dots,z_{n}$ in the integral basis $(1,\omega_{D})$. We call $Q$ the quadratic form associated to $G$. It has coefficients in $K$ and the variables $x_{1},y_{1},\dots,x_{n},y_{n}$ take values in $\mathbb{Q}$.
@@
for $\mathbf{z}\in K^{n}$.
@@
determines the change of variables $(x_{1},y_{1},\dots,x_{n},y_{n})\mapsto(z_{1},\dots,z_{n},\tau(z_{1}),\dots,\tau(z_{n}))$. Thus, if $M_{Q}$ is the matrix of the associated quadratic form $Q$ defined by (4), then
@@
$\det(M_{Q})=D^{n}\cdot\begin{cases}\det(2M_{G}),&\text{if }D\equiv 2,3\pmod{4},\\ \det(M_{G}),&\text{if }D\equiv 1\pmod{4}.\end{cases}$
@@
$\det T=(-1)^{\frac{n(n-1)}{2}}(\det M)^{n}=\begin{cases}(-1)^{\frac{n(n+1)}{2}}D^{\frac{n}{2}}2^{n},&\text{if }D\equiv 2,3\pmod{4},\\ (-1)^{\frac{n(n+1)}{2}}D^{\frac{n}{2}},&\text{if }D\equiv 1\pmod{4}.\end{cases}$
@@
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree and let $G$ be a generalized form in $n$ variables given by (3). Assume that $G(\mathbf{a})\in\mathbb{Q}$ for every $\mathbf{a}\in K^{2}$. If $1\leq i\leq j\leq n$, then we let $\alpha_{ij}=r_{ij}+s_{ij}\omega_{D}$ and $\beta_{ij}=t_{ij}+u_{ij}\omega_{D}$ where $r_{ij},s_{ij},t_{ij},u_{ij}\in\mathbb{Q}$. Let $Q$ be the quadratic form in $2n$ variables associated to $G$. If $1\leq i\leq j\leq n$, then we let $a_{ij}$, $b_{ij}$, $b_{ji}$, and $c_{ij}$ be the coefficients of the terms $x_{i}x_{j}$, $x_{i}y_{j}$, $x_{j}y_{i}$, and $y_{i}y_{j}$ in $Q$, respectively.
@@
We introduce the following notation: if $M$ is a $2n\times 2n$ matrix and $1\leq k\leq 2n$, let $[M]_{k}$ denote the $k$-th row of $M$.
@@
$[M_{Q}]_{2k}=\begin{pmatrix}\frac{b_{1k}}{2}&\frac{c_{1k}}{2}&\dots&\frac{b_{kk}}{2}&c_{kk}&\dots&\frac{b_{nk}}{2}&\frac{c_{kn}}{2}\end{pmatrix}.$
@@
Since we assume that $G$ is integral, $r_{ij},s_{ij},t_{ij},u_{ij}\in\mathbb{Z}$. Thus $D$ divides every element of $[M_{Q}]_{2k}$. It follows that $M_{Q}\bmod{p}$ contains $n$ zero rows, hence $\operatorname{rank}(M_{Q}\bmod{p})\leq n$.
Secondly, assume that $D\equiv 1\pmod{4}$. Let $N_{Q}$ be the matrix obtained from $M_{Q}$ by subtracting two times the $2k$-th row of $M_{Q}$ from the $(2k-1)$-th row of $M_{Q}$ for every $1\leq k\leq n$. The $(2k-1)$-th row of $N_{Q}$ equals
$[M_{Q}]_{2k-1}-2[M_{Q}]_{2k}=\begin{pmatrix}\frac{a_{1k}}{2}-b_{1k}&\frac{b_{k1}}{2}-c_{1k}&\dots&a_{kk}-b_{kk}&\frac{b_{kk}}{2}-2c_{kk}&\dots&\frac{a_{kn}}{2}-b_{nk}&\frac{b_{kn}}{2}-c_{kn}\end{pmatrix}.$
@@
We have $r_{ij},s_{ij},t_{ij},u_{ij}\in\frac{1}{2}\mathbb{Z}$ because $G$ is integral. Since we assume $D\equiv 1\pmod{4}$, $p$ is an odd prime and $2$ is invertible modulo $p$. We see that every element of $[N_{Q}]_{2k-1}$ vanishes modulo $p$, hence $N_{Q}\bmod{p}$ contains $n$ zero rows and $\operatorname{rank}(M_{Q}\bmod{p})=\operatorname{rank}(N_{Q}\bmod{p})\leq n$. ∎
@@ 5. Binary and ternary generalized quadratic forms
In Lemmas 5.4 to 5.9, we find an example of a $\mathbb{Z}$-universal generalized form over $\mathbb{Q}(\sqrt{D})$ for $D\in\{2,3,6,7,10\}$ and of a classical $\mathbb{Z}$-universal generalized form over $\mathbb{Q}(\sqrt{5})$.
@@
The coefficients of $G_{2}$ are $a=1$, $b=-\sqrt{2}$, $c=1$, $d=-1$, $e=0$, $f=-1$. Let $Q_{2}$ be the quaternary quadratic form $Q_{2}(x_{1},y_{1},x_{2},y_{2})=G_{2}(x_{1}+y_{1}\sqrt{2},x_{2}+y_{2}\sqrt{2})$ and let $M_{Q_{2}}$ be the matrix of $Q_{2}$. By Lemma 5.3 with $D=2$,
@@
Proof.
The coefficients of $G_{3}$ are $a=2$, $b=-2\sqrt{3}$, $c=2$, $d=-3$, $e=-\sqrt{3}$, $f=0$. Let $Q_{3}$ be the quaternary quadratic form $Q_{3}(x_{1},y_{1},x_{2},y_{2})=G_{3}(x_{1}+y_{1}\sqrt{3},x_{2}+y_{2}\sqrt{3})$. By Lemma 5.3 with $D=3$,
$M_{Q_{3}}=\begin{pmatrix}2a_{1}+d&6a_{2}&b_{1}+e_{1}&3(b_{2}-e_{2})\\ 6a_{2}&3(2a_{1}-d)&3(b_{2}+e_{2})&3(b_{1}-e_{1})\\ b_{1}+e_{1}&3(b_{2}-e_{2})&2c_{1}+f&6c_{2}\\ 3(b_{2}-e_{2})&3(b_{1}-e_{1})&6c_{2}&3(2c_{1}-f)\end{pmatrix}=\begin{pmatrix}1&0&0&-3\\ 0&21&-9&0\\ 0&-9&4&0\\ -3&0&0&12\end{pmatrix}.$
We get
$M_{Q_{3}}\sim\begin{pmatrix}1&0&0&0\\ 0&21&-9&0\\ 0&-9&4&0\\ 0&0&0&3\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&1&-1&0\\ 0&-1&4&0\\ 0&0&0&3\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&3&0\\ 0&0&0&3\end{pmatrix},$
The diagonal quadratic form $x^{2}+y^{2}+3z^{2}+3w^{2}$ is positive definite and universal by the 15-Theorem, hence $G_{3}$ is also positive definite and $\mathbb{Z}$-universal. ∎
@@
Proof.
The coefficients of $G_{5}$ are $a=3-\omega_{5}$, $b=0$, $c=1$, $d=-4$, $e=0$, and $f=0$. Since the coefficients of the off-diagonal terms are even, $G_{5}$ is classical. Let $Q_{5}$ be the quaternary quadratic form $Q_{5}(x_{1},y_{1},x_{2},y_{2})=G_{5}(x_{1}+y_{1}\omega_{5},x_{2}+y_{2}\omega_{5})$. By Lemma 5.3 with $D=5$,
$M_{Q_{5}}=\begin{pmatrix}M_{Q_{5},1}&M_{Q_{5},2}\\ M_{Q_{5},2}^{\top}&M_{Q_{5},3}\end{pmatrix}$
where
$M_{Q_{5},1}$ $=\begin{pmatrix}2a_{1}+a_{2}+d&a_{1}+3a_{2}+\frac{d}{2}\\ a_{1}+3a_{2}+\frac{d}{2}&3a_{1}+4a_{2}-d\end{pmatrix}=\begin{pmatrix}1&-2\\ -2&9\end{pmatrix},$
$M_{Q_{5},2}$ $=\begin{pmatrix}b_{1}+\frac{b_{2}}{2}+e_{1}+\frac{e_{2}}{2}&\frac{b_{1}}{2}+\frac{3}{2}b_{2}+\frac{e_{1}}{2}-e_{2}\\ \frac{b_{1}}{2}+\frac{3}{2}b_{2}+\frac{e_{1}}{2}+\frac{3}{2}e_{2}&\frac{3}{2}b_{1}+2b_{2}-e_{1}-\frac{1}{2}e_{2}\end{pmatrix}=\begin{pmatrix}0&0\\ 0&0\end{pmatrix},$
$M_{Q_{5},3}$ $=\begin{pmatrix}2c_{1}+c_{2}+f&c_{1}+3c_{2}+\frac{f}{2}\\ c_{1}+3c_{2}+\frac{f}{2}&3c_{1}+4c_{2}-f\end{pmatrix}=\begin{pmatrix}2&1\\ 1&3\end{pmatrix},$
thus
$M_{Q_{5}}=\begin{pmatrix}1&-2&0&0\\ -2&9&0&0\\ 0&0&2&1\\ 0&0&1&3\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&5&0&0\\ 0&0&2&1\\ 0&0&1&3\end{pmatrix}.$
This quadratic form is positive definite and universal by the 15-Theorem. Alternatively, it is possible to consult [Bh, Table 5], which contains the equivalent form
$\begin{pmatrix}1&0&0&0\\ 0&2&1&0\\ 0&1&3&0\\ 0&0&0&5\end{pmatrix}$
encoded as “25: 2 3 5 0 0 2:2523500225:\ 2\ 3\ 5\ 0\ 0\ 225 : 2 3 5 0 0 2”. ∎
@@
Proof.
The coefficients of $G_{6}$ are $a=3$, $b=2\sqrt{6}$, $c=3$, $d=-5$, $e=\sqrt{6}$, and $f=-1$. Let $Q_{6}$ be the quaternary quadratic form $M_{Q_{6}}(x_{1},y_{1},x_{2},y_{2})=G_{6}(x_{1}+y_{1}\sqrt{6},x_{2}+y_{2}\sqrt{6})$. By Lemma 5.3 with $D=6$,
$M_{Q_{6}}=\begin{pmatrix}2a_{1}+d&12a_{2}&b_{1}+e_{1}&6(b_{2}-e_{2})\\ 12a_{2}&6(2a_{1}-d)&6(b_{2}+e_{2})&6(b_{1}-e_{1})\\ b_{1}+e_{1}&6(b_{2}+e_{2})&2c_{1}+f&12c_{2}\\ 6(b_{2}-e_{2})&6(b_{1}-e_{1})&12c_{2}&6(2c_{1}-f)\end{pmatrix}=\begin{pmatrix}1&0&0&6\\ 0&66&18&0\\ 0&18&5&0\\ 6&0&0&42\end{pmatrix}.$
We get
$M_{Q_{6}}\sim\begin{pmatrix}1&0&0&0\\ 0&66&18&0\\ 0&18&5&0\\ 0&0&0&6\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&3&3&0\\ 0&3&5&0\\ 0&0&0&6\end{pmatrix}\sim\begin{pmatrix}1&0&0&0\\ 0&3&0&0\\ 0&0&2&0\\ 0&0&0&6\end{pmatrix}.$
The last quadratic form is positive definite and universal by the 15-Theorem. ∎
@@
Proof.
The coefficients of $G_{7}$ are $a=3$, $b=1+2\sqrt{7}$, $c=4$, $d=-5$, $e=-1+2\sqrt{7}$, and $f=3$. By Lemma 5.3 with $D=7$,
$M_{Q_{7}}=\begin{pmatrix}2a_{1}+d&14a_{2}&b_{1}+e_{1}&7(b_{2}-e_{2})\\ 14a_{2}&7(2a_{1}-d)&7(b_{2}+e_{2})&7(b_{1}-e_{1})\\ b_{1}+e_{1}&7(b_{2}+e_{2})&2c_{1}+f&14c_{2}\\ 7(b_{2}-e_{2})&7(b_{1}-e_{1})&14c_{2}&7(2c_{1}-f)\end{pmatrix}=\begin{pmatrix}1&0&0&0\\ 0&77&28&14\\ 0&28&11&0\\ 0&14&0&35\end{pmatrix}.$
Let
$A_{7}=\begin{pmatrix}1&0&0&0\\ 0&2&-3&3\\ 0&-5&8&-7\\ 0&-1&1&-1\end{pmatrix}.$
It can be verified that $\det(A_{7})=1$, hence $A_{7}\in\operatorname{SL}_{4}(\mathbb{Z})$. The equivalent form with the matrix
$A_{7}^{\top}M_{Q_{7}}A_{7}=\begin{pmatrix}1&0&0&0\\ 0&2&-3&3\\ 0&-5&8&-7\\ 0&-1&1&-1\end{pmatrix}^{\top}\begin{pmatrix}1&0&0&0\\ 0&77&28&14\\ 0&28&11&0\\ 0&14&0&35\end{pmatrix}\begin{pmatrix}1&0&0&0\\ 0&2&-3&3\\ 0&-5&8&-7\\ 0&-1&1&-1\end{pmatrix}=\begin{pmatrix}1&0&0&0\\ 0&2&1&0\\ 0&1&4&0\\ 0&0&0&7\end{pmatrix}$
is positive definite and universal by the 15-Theorem. Alternatively, it can be found in [Bh, Table 5] as “49: 2 4 7 0 0 2”. ∎
@@
Proof.
The coefficients of $G_{10}$ are $a=3$, $b=\sqrt{10}$, $c=5$, $d=-5$, $e=\sqrt{10}$, and $f=7$. Let $Q_{10}$ be the quaternary quadratic form $Q_{10}(x_{1},y_{1},x_{2},y_{2})=G_{10}(x_{1}+y_{1}\sqrt{10},x_{2}+y_{2}\sqrt{10})$. By Lemma 5.3 with $D=10$,
$M_{Q_{10}}=\begin{pmatrix}2a_{1}+d&20a_{2}&b_{1}+e_{1}&10(b_{2}-e_{2})\\ 20a_{2}&10(2a_{1}-d)&10(b_{2}+e_{2})&10(b_{1}-e_{1})\\ b_{1}+e_{1}&10(b_{2}+e_{2})&2c_{1}+f&20c_{2}\\ 10(b_{2}-e_{2})&10(b_{1}-e_{1})&20c_{2}&10(2c_{1}-f)\end{pmatrix}=\begin{pmatrix}1&0&0&0\\ 0&110&20&0\\ 0&20&17&20\\ 0&0&20&30\end{pmatrix}.$
Let
$A_{10}=\begin{pmatrix}1&0&0&0\\ 0&1&1&2\\ 0&-6&-5&-10\\ 0&4&3&7\end{pmatrix}.$
It can be verified that $\det(A_{10})=1$, hence $A\in\operatorname{SL}_{4}(\mathbb{Z})$. The equivalent form with the matrix
$A_{10}^{\top}M_{Q_{10}}A_{10}=\begin{pmatrix}1&0&0&0\\ 0&1&1&2\\ 0&-6&-5&-10\\ 0&4&3&7\end{pmatrix}^{\top}\begin{pmatrix}1&0&0&0\\ 0&110&20&0\\ 0&20&17&20\\ 0&0&20&30\end{pmatrix}\begin{pmatrix}1&0&0&0\\ 0&1&1&2\\ 0&-6&-5&-10\\ 0&4&3&7\end{pmatrix}=\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&5&0\\ 0&0&0&10\end{pmatrix}$
is positive definite and universal by the 15-Theorem. ∎
@@
First we prove (i). We assume that $D\equiv 2,3\pmod{4}$. If a binary $\mathbb{Z}$-universal generalized form exists over $K=\mathbb{Q}(\sqrt{D})$, then $D\in\{2,3,6,7,10\}$ by Proposition 5.2 (i). Conversely, in Lemmas 5.4, 5.5, 5.7, 5.8 and 5.9 we found an example of a generalized form with the required properties over each of these fields.
@@
This proves $D\mid\det(M_{R})$. Every quaternary form obtained by the method of escalations is listed in [Bh, Table 3] and in particular, has determinant $\leq 112$. It remains to determine which squarefree $D\leq 112$ divide the determinant of some form on the list. Every squarefree $D\in\{1,\dots,62\}$ satisfies this condition. Additionally, the squarefree determinants greater than $62$ on the list are $66,70,74,77,78,82,85,86,87,93,94,95,102,106,110$. ∎
@@
We show that there exists a positive constant $C$ such that $D\leq C$ for every such field $K=\mathbb{Q}(\sqrt{D})$. If $D\equiv 2,3\pmod{4}$, then the statement follows from Theorems 5.10 and 5.11. Hence, we can assume that $D\equiv 1\pmod{4}$. Let $G$ be a $\mathbb{Z}$-universal generalized form over $K$ with matrix $M_{G}$ and $Q$ be the quadratic form associated to $G$ with matrix $M_{Q}$.
@@
We end this section with a proposition that will be used in Section 7.
@@
$\det(2M_{G})=2a\begin{vmatrix}2c&\tau(e)&f\\ \tau(e)&2\tau(a)&\tau(b)\\ f&\tau(b)&2\tau(c)\end{vmatrix}-b\begin{vmatrix}b&d&e\\ \tau(e)&2\tau(a)&\tau(b)\\ f&\tau(b)&2\tau(c)\end{vmatrix}+d\begin{vmatrix}b&d&e\\ 2c&\tau(e)&f\\ f&\tau(b)&2\tau(c)\end{vmatrix}-e\begin{vmatrix}b&d&e\\ 2c&\tau(e)&f\\ \tau(e)&2\tau(a)&\tau(b)\end{vmatrix},$
@@
$\det(2M_{G})\equiv$ $2a\left[\tau(e)\tau(b)f+f\tau(e)\tau(b)\right]-b\left[d\tau(b)f+e\tau(e)\tau(b)-2f\tau(a)e-\tau(b)^{2}b-2\tau(c)\tau(e)d\right]$
$+d\left[2b\tau(e)\tau(c)+df^{2}+2ec\tau(b)-f\tau(e)e-\tau(b)fb\right]$
$-e\left[b\tau(e)\tau(b)+df\tau(e)-\tau(e)^{2}e-2\tau(a)fb-2\tau(b)cd\right]$
@@
$\frac{\det(M_{Q})}{D^{2}}=\det(2M_{G})\equiv 0,1\pmod{4}.\qed$

### 21 Mar 2026 v2

@@ 3. Generalized quadratic forms
Let us recall that a quadratic form $Q$ in $n$ variables over $K$ is a homogeneous polynomial of degree $2$ over $K$. We can also obtain it as a generalized quadratic form with $\beta_{ij},\gamma_{ij}=0$ for all $1\leq i,j\leq n.$
Then if $G$ is a generalized quadratic form in $n$ variables over $K$, then we can define a quadratic form $Q$ associated to $G$ in $2n$ variables by
@@
In other words, $Q$ is obtained from $G$ by expressing the variables $z_{1},z_{2},\dots,z_{n}$ in the integral basis $(1,\omega_{D})$. It has coefficients in $K$ and the variables $x_{1},y_{1},\dots,x_{n},y_{n}$ take values in $\mathbb{Q}$.
@@
for $\mathbf{z}\in K^{n}$. The rank of $G$ is the rank of the matrix $M_{G}$ over $K$.
@@
determines the change of variables $(x_{1},y_{1},\dots,x_{n},y_{n})\mapsto(z_{1},\dots,z_{n},\tau(z_{1}),\dots,\tau(z_{n}))$. Thus, if is the quadratic form $Q$ is associated to the quadratic form $G$ defined by (4), and $M_{Q},M_{G}$ are their respective matrices, then
@@
If $G$ is a generalized quadratic form over $\mathbb{Q}$ then we say that $G$ represents $b\in\mathbb{Z}$ over $\mathbb{Z}$ if there exists $\mathbf{a}\in\mathbb{Z}^{n}$ such that $G(\mathbf{a})=b$. In particular, according to our definition, every quadratic form represents 0. If there exists $\mathbf{a}\in\mathbb{Z}^{n}\setminus\{(0,0,\dots,0)\}$ such that $G(\mathbf{a})=0$, then we say that $G$ represents 0 non-trivially. The form $G$ is called positive definite if $M_{G}$ is a positive definite matrix and $G$ is called universal if it is a positive definite integral form such that it represents every $b\in\mathbb{Z}_{\geq 1}$.
@@
[…]
@@
[…]
@@
Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree and let $G$ be a generalized form in $n$ variables given by (3). Assume that $G(\mathbf{a})\in\mathbb{Q}$ for every $\mathbf{a}\in K^{n}$. If $1\leq i\leq j\leq n$, then we let $\alpha_{ij}=r_{ij}+s_{ij}\omega_{D}$ and $\beta_{ij}=t_{ij}+u_{ij}\omega_{D}$ where $r_{ij},s_{ij},t_{ij},u_{ij}\in\mathbb{Q}$. Let $Q$ be the quadratic form in $2n$ variables associated to $G$. If $1\leq i\leq j\leq n$, then we let $a_{ij}$, $b_{ij}$, $b_{ji}$, and $c_{ij}$ be the coefficients of the terms $x_{i}x_{j}$, $x_{i}y_{j}$, $x_{j}y_{i}$, and $y_{i}y_{j}$ in $Q$, respectively.
@@
We introduce the following notation: if $M$ is a $2n\times 2n$ matrix and $1\leq k\leq 2n$, let $M]_{k}$ denote the $k$-th row of $M$.
@@
$M_{Q}]_{2k}=\begin{pmatrix}\frac{b_{1k}}{2}&\frac{c_{1k}}{2}&\dots&\frac{b_{kk}}{2}&c_{kk}&\dots&\frac{b_{nk}}{2}&\frac{c_{kn}}{2}\end{pmatrix}$.
@@
Since we assume that $G$ is integral, $r_{ij},s_{ij},t_{ij},u_{ij}\in\mathbb{Z}$. Thus $D$ divides every element of $M_{Q}]_{2k}$. It follows that $M_{Q}\bmod{p}$ contains $n$ zero rows, hence $\operatorname{rank}(M_{Q}\bmod{p})\leq n$.
Secondly, assume that $D\equiv 1\pmod{4}$. Let $N_{Q}$ be the matrix obtained from $M_{Q}$ by subtracting two times the $2k$-th row of $M_{Q}$ from the (2k1)(2k-1)-th row of $M_{Q}$ for every $1\leq k\leq n$. The (2k1)(2k-1)-th row of $N_{Q}$ equals
$M_{Q}]_{2k-1}-2[M_{Q}]_{2k}=\begin{pmatrix}\frac{a_{1k}}{2}-b_{1k}&\frac{b_{k1}}{2}-c_{1k}&\dots&a_{kk}-b_{kk}&\frac{b_{kk}}{2}-2c_{kk}&\dots&\frac{a_{kn}}{2}-b_{nk}&\frac{b_{kn}}{2}-c_{kn}\end{pmatrix}$.
@@
We have $r_{ij},s_{ij},t_{ij},u_{ij}\in\frac{1}{2}\mathbb{Z}$ because $G$ is integral. Since we assume $D\equiv 1\pmod{4}$, $p$ is an odd prime and $2$ is invertible modulo $p$. We see that every element of $N_{Q}]_{2k-1}$ vanishes modulo $p$, hence $N_{Q}\bmod{p}$ contains $n$ zero rows and $\operatorname{rank}(M_{Q}\bmod{p})=\operatorname{rank}(N_{Q}\bmod{p})\leq n$. ∎
@@ 4. Binary and ternary generalized quadratic forms
In Lemmas 4.4 to 4.9, we find an example of a $\mathbb{Z}$-universal generalized form over $\mathbb{Q}(\sqrt{D})$ for $D\in\{2,3,6,7,10\}$ and of a classical $\mathbb{Z}$-universal generalized form over $\mathbb{Q}(\sqrt{5})$. We prove only Lemma 4.4, the rest is stated and proved in the Appendix in Section A4.
@@
The coefficients of $G_{2}$ are $a=1$, $b=-\sqrt{2}$, $c=1$, d=1d=-1, $e=0$, f=1f=-1. Let $Q_{2}$ be the quaternary quadratic form $Q_{2}(x_{1},y_{1},x_{2},y_{2})=G_{2}(x_{1}+y_{1}\sqrt{2},x_{2}+y_{2}\sqrt{2})$ and let $M_{Q_{2}}$ be the matrix of $Q_{2}$. By Lemma 4.3 with $D=2$,
@@
First we prove (i). We assume that $D\equiv 2,3\pmod{4}$. If a binary $\mathbb{Z}$-universal generalized form exists over $K=\mathbb{Q}(\sqrt{D})$, then $D\in\{2,3,6,7,10\}$ by Proposition 4.2 (i). Conversely, in Lemmas 4.4, 4.5, 4.7, 4.8 and 4.9 we found an example of a generalized form with the required properties over each of these fields.
@@
This proves $D\mid\det(M_{R})$. Every quaternary form obtained by the method of escalations is listed in [Bh, Table 3] and in particular, has determinant $\leq 112$. It remains to determine which squarefree $D\leq 112$ divide the determinant of some form on the list. Every squarefree $D\in\{1,\dots,62\}$ satisfies this condition. Additionally, the squarefree determinants greater than $62$ on the list are 66,70,74,77,78,82,85,86,87,93,94,95,102,106,11066,70,74,77,78,82,85,86,87,93,94,95,102,106,110. ∎
@@
We show that there exists a positive constant $C$ such that $D\leq C$ for every such field $K=\mathbb{Q}(\sqrt{D})$. If $D\equiv 2,3\pmod{4}$, then the statement follows from Theorems 4.10 and 4.11. Hence, we can assume that $D\equiv 1\pmod{4}$. Let $G$ be a $\mathbb{Z}$-universal generalized form over $K$ with matrix $M_{G}$ and $Q$ be the quadratic form associated to $G$ with matrix $M_{Q}$.
@@
We end this section with a proposition that is used in Section A6.
@@
$\det(2M_{G})=$ $2a\cdot\det\begin{pmatrix}2c&\tau(e)&f\\ \tau(e)&2\tau(a)&\tau(b)\\ f&\tau(b)&2\tau(c)\end{pmatrix}-b\cdot\det\begin{pmatrix}b&d&e\\ \tau(e)&2\tau(a)&\tau(b)\\ f&\tau(b)&2\tau(c)\end{pmatrix}$+
$+d\cdot\det\begin{pmatrix}b&d&e\\ 2c&\tau(e)&f\\ f&\tau(b)&2\tau(c)\end{pmatrix}-e\cdot\det\begin{pmatrix}b&d&e\\ 2c&\tau(e)&f\\ \tau(e)&2\tau(a)&\tau(b)\end{pmatrix}$,
@@
$\det(2M_{G})\equiv$ $2a\Big(\tau(e)\tau(b)f+f\tau(e)\tau(b)\Big)-b\Big(d\tau(b)f+e\tau(e)\tau(b)-2f\tau(a)e-\tau(b)^{2}b-2\tau(c)\tau(e)d\Big)$
$+d\Big(2b\tau(e)\tau(c)+df^{2}+2ec\tau(b)-f\tau(e)e-\tau(b)fb\Big)$
$-e\Big(b\tau(e)\tau(b)+df\tau(e)-\tau(e)^{2}e-2\tau(a)fb-2\tau(b)cd\Big)$
@@
$\frac{\det(M_{Q})}{D^{2}}=\det(2M_{G})\equiv 0,1\pmod{4}$$\qed$